# Online Multi-Agent Pickup and Delivery with Task Deadlines

### v1

@@ V-A Evaluation of Task Tardiness
The numerical experiments were carried out in a grid environment, representing an automated warehouse, as illustrated in Fig. 1. We generated 151 tasks by randomly selecting pickup and delivery locations from the task endpoints, ensuring no duplication. Each agent moved to the delivery location via the pickup location for the assigned task. We varied parameters such as task-release times and deadline duration after task release to examine their significant impact during the experiments (see Table II). The experiment involved 15 agents, with their initial positions randomly selected from the non-task endpoints.
@@
*in [min, max].
Figure 2: Comparisons of cumulative tardines.
We also examined an offline setting where tasks and their release times are predetermined, in contrast to the online setting where tasks can be added at any time. In an offline setting, tasks can be allocated with foresight to accommodate future tasks and preemptively moved in anticipation of release times. Generally, offline methods tend to yield solutions closer to the optimal solution compared to online methods. In this context, we referred to the method by Ramanathan et al. [15], who regarded deadlines in offline tasks as ideal benchmarks. They employed a method that sorted tasks based on deadlines and assigned tasks to agents with lower execution costs.
Fig. 2 illustrates the results of the proposed method alongside ideal values in the offline setting. The horizontal axis represents the weight $\alpha$ used during task assignment, where smaller values prioritize execution cost and larger values prioritize deadlines. When $\alpha=0$ and task switching is not employed, the method is equivalent to TP and TPTS [9]. The vertical axis indicates the total tardiness for each task, averaged across the results of 30 experiments. Our analysis demonstrates that the cumulative tardiness varies depending on the value of $\alpha$ and the presence of task exchanges, regardless of the release frequency and deadlines.
@@ V-B Discussion
Next, we explore the weight $\alpha$ in the proposed method. We compare the disparities in cumulative tardiness caused by varying $\alpha$ when neither task swapping nor task switching is implemented. Minimum tardiness is observed for dense task releases and short deadlines (Fig. 2(a)) at $\alpha=0.0$, for dense releases and long deadlines (Fig. 2(b)) at $\alpha=0.1$, for sparse releases and short deadlines (Fig. 2(c)) at $\alpha=0.025$, and for sparse releases and long deadlines (Fig. 2(d)) at $\alpha=0.2$. When tasks are released frequently and deadlines are short, prioritizing tasks with lower execution costs is crucial for timely management. However, in scenarios where tasks are not released frequently or deadlines are not short, there is some flexibility to consider deadlines. While TP and TPTS [9] solely considered execution costs ($\alpha=0$), the proposed method integrates temporal margin for deadlines along with execution costs, resulting in reduced tardiness. Nonetheless, excessively prioritizing the temporal margin of deadlines may escalate the execution costs of each task. This can overwhelm agents and increasing cumulative tardiness. Hence, adjusting the value of $\alpha$ based on the situation is imperative.
@@
In all experiments, the lowest tardiness was achieved when both task swapping and task switching were implemented. For example, in Fig. 2(b), implementing only task swapping reduced cumulative tardiness by $569.4$, whereas implementing both task swapping and task switching further reduced tardiness by an additional $116.7$. Task switching facilitates task reassignment when more urgent tasks are added. Additionally, task swapping enables task exchanges between agents, leading to further reduction in cumulative tardiness.
@@
However, implementing only task switching may elevate cumulative tardiness compared to not implementing anything, especially when task releases are frequent (Figs. 2(a) and 2(b)). This aligns with the discussion on the value of $\alpha$; in scenarios with frequent task releases, minimizing execution costs outweighs considering task urgency. Task reassignment based on urgency through task switching increases execution costs, consequently amplifying tardiness.
@@
Finally, we compare the proposed method (online) with the ideal values (offline). In most cases, the tardiness of the proposed method was equivalent to or worse than the ideal values (Figs. 2(a), (c), and (d)). However, in scenarios with dense tasks and long deadlines, the proposed method outperformed the ideal values (Fig. 2(b)). Ramanathan et al. [15] sorted tasks in advance based on deadlines and assigned them to agents with lower execution costs. This indicated that they prioritized deadlines over execution costs. They noted that their method excelled with extremely short deadlines, exhibiting less tardiness than the proposed method in settings with frequent releases and short deadlines. However, maintaining a balance between execution costs and tardiness becomes crucial when handling numerous tasks and longer deadlines. The effectiveness of the proposed method is evidenced in such scenarios despite operating online.
@@ VI Conclusions
This study addresses task deadlines by introducing a modified version of the MAPD problem, termed online MAPD-D. In online MAPD-D, tasks can be added at any time and assigned deadlines. To address MAPD-D, we propose two algorithms, namely D-TP and D-TPTS. D-TP allocates tasks by considering their pickup deadlines along with execution costs; meanwhile, D-TPTS facilitates task exchanges among agents and within a single agent. The conducted numerical experiments demonstrate that both D-TP and D-TPTS effectively reduce task tardiness compared to existing methods.
These experiments are conducted in a $35\times 21$ grid environment; however, our ability to solve MAPD-D in larger environments is limited due to computational constraints. In the future, exploring the development of decentralized algorithms could enable the solution of large-scale MAPD-D. Additionally, algorithms should be devised to handle more realistic scenarios, such as paths being obstructed by uncertain obstacles [19].

### v2

@@ V-A Evaluation of Task Tardiness
Numerical experiments were conducted in a grid environment, representing an automated warehouse, as illustrated in Fig. 3. We generated 151 tasks by selecting pickup and delivery locations uniformly at random from the task endpoints, ensuring no duplication. Each agent moved to the delivery location via the pickup location for the assigned task. We varied parameters such as task-release times and deadline duration after task release to examine their significant impact during the experiments (see Table II). The experiment involved 15 agents, with their initial positions randomly selected from the non-task endpoints.
Figure 3: Four-neighbor grid environment [9]. A pair of pickup and delivery locations with the same color represents a single task. In MAPD-D, each delivery location is assigned a deadline.
We also examined an offline setting where tasks and their release times are predetermined, contrary to the online setting where tasks can be added at any time. In an offline setting, tasks can be allocated with foresight to accommodate future tasks and preemptively moved in anticipation of release times. Generally, offline methods tend to yield solutions closer to the optimal solution compared to online methods. Therefore, we referred to the method by Ramanathan et al. [15], who regarded deadlines in offline tasks as ideal benchmarks. They employed a method that sorted tasks based on deadlines and assigned tasks to agents with lower execution costs.
Fig. 4 illustrates the results of the proposed method alongside ideal values in the offline setting. The horizontal axis represents the weight $\alpha$ used during task assignment, where smaller values prioritize execution cost and larger values prioritize deadlines. When $\alpha=0$ and task switching is not employed, our proposed new task swapping method is equivalent to TP and TPTS [9]. The vertical axis indicates the total tardiness for each task, averaged across the results of 30 experiments. Our analysis demonstrates that the cumulative tardiness varies depending on the value of $\alpha$ and the presence of task exchanges, regardless of the release frequency and deadlines.
@@
in [min, max].
Figure 4: Comparisons of cumulative tardiness.
@@ V-B Discussion
Next, we explore the weight $\alpha$ in the proposed method. We compare the disparities in cumulative tardiness caused by varying $\alpha$ when neither task swapping nor task switching is implemented. Minimum tardiness is observed for dense task releases and short deadlines (Fig. 4(a)) at $\alpha=0.0$, for dense releases and long deadlines (Fig. 4(b)) at $\alpha=0.1$, for sparse releases and short deadlines (Fig. 4(c)) at $\alpha=0.025$, and for sparse releases and long deadlines (Fig. 4(d)) at $\alpha=0.2$. When tasks are released frequently and deadlines are short, prioritizing tasks with lower execution costs is crucial for timely management. However, in scenarios with less frequent task releases or longer deadlines, there is some flexibility to consider deadlines. While TP and TPTS [9] solely considered execution costs ($\alpha=0$), the proposed method integrates temporal margin for deadlines along with execution costs, resulting in reduced tardiness. Nonetheless, excessively prioritizing the temporal margin of deadlines may escalate the execution costs of each task. This can overwhelm agents and increase cumulative tardiness. Hence, adjusting the value of $\alpha$ based on the situation is imperative.
@@
In all experiments, the lowest tardiness was achieved when both task swapping and task switching were implemented. For example, as shown in Fig. 4(b), implementing only task swapping reduced cumulative tardiness by $569.4$, whereas implementing both task swapping and task switching further reduced tardiness by an additional $116.7$. Task switching facilitates task reassignment when more urgent tasks are added. Additionally, task swapping enables task exchanges between agents, further reducing cumulative tardiness.
@@
However, implementing only task switching may elevate cumulative tardiness compared to not implementing any strategy, particularly when task releases are frequent (Figures 4(a) and 4(b)). This aligns with the discussion on the value of $\alpha$; in scenarios with frequent task releases, minimizing execution costs outweighs considering task urgency. Task reassignment based on urgency through task switching increases execution costs, consequently amplifying tardiness.
@@
Finally, we compare the proposed method (online) with the ideal values (offline). In most cases, the tardiness of the proposed method was equivalent to or worse than the ideal values (Figures 4(a), (c), and (d)). However, in scenarios with dense tasks and long deadlines, the proposed method outperformed the ideal values (Fig. 4(b)). Ramanathan et al. [15] sorted tasks in advance based on deadlines and assigned them to agents with lower execution costs. This indicates that they prioritized deadlines over execution costs. They noted that their method excelled with extremely short deadlines, exhibiting less tardiness than the proposed method in settings with frequent releases and short deadlines. However, maintaining a balance between execution costs and tardiness becomes crucial when handling numerous tasks and longer deadlines. The effectiveness of the proposed method is evidenced in such scenarios despite operating online.
@@ VI Conclusions
This study addresses task deadlines by introducing a modified version of the MAPD problem, termed online MAPD-D. In online MAPD-D, tasks can be added at any time and assigned deadlines. To address MAPD-D, we propose two algorithms: D-TP and D-TPTS. D-TP allocates tasks by considering their pickup deadlines along with execution costs, while D-TPTS facilitates task exchanges among agents and within a single agent. The numerical experiments demonstrated that both D-TP and D-TPTS effectively reduced task tardiness compared with existing methods.
These experiments were conducted in a $35\times 21$ grid environment; however, our ability to solve MAPD-D in larger environments is limited owing to computational constraints. In the future, exploring the development of decentralized algorithms could enable the solution of large-scale MAPD-D. Additionally, algorithms should be devised to handle more realistic scenarios, such as paths being obstructed by uncertain obstacles [22]. Furthermore, optimizing the parameter $\alpha$ in (3) is necessary. While $\alpha$ can be tuned based on past trends, it is challenging to tune $\alpha$ for future trends in online task settings.